# DOPRA: Decoding Over-accumulation Penalization and Re-allocation in Specific Weighting Layer

### v1

@@ Abstract.
* represents the same contribution, Xiaofeng Zhang is the corresponding author, his email:framebreak@sjtu.edu.cn
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 28 - November 1, 2024; Melbourne, Australia.isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Recently, Multimodal Large Language Models (MLLMs)(Achiam et al., 2023; Liu et al., 2024; Bai et al., 2023; Zhu et al., 2023; Dai et al., 2024; Li et al., 2023d; Team et al., 2023; Gong et al., 2023; Zhang et al., 2023a) have made groundbreaking advancements, fundamentally altering the way AI interacts with visual data and significantly enhancing fluent communication based on image semantic content. Despite their remarkable performance in handling a range of visually-centered tasks (Zhang et al., 2023c; Black et al., 2023; Li et al., 2024; Chen et al., 2019; Cai et al., 2023), understanding complex contexts(Lai et al., 2023; Zhao et al., 2023a), or generating coherent narratives (Brooks et al., 2023; Black et al., 2023; Li et al., 2024), MLLMs still grapple with a profound challenge: the ”hallucination” problem. This refers to instances where MLLMs generate inaccurate or disjointed responses to visual inputs by incorrectly identifying nonexistent objects, attributes, or relationships within provided images. Such errors carry significant risks, particularly in high-stakes applications like autonomous driving (Muhammad et al., 2020; Shafaei et al., 2018), where misinterpreting visual cues could lead to life-threatening situations. While numerous methods (Liu et al., 2023a; Park et al., 2024; Yin et al., 2023; Wang et al., 2024; Zhou et al., 2023) have been proposed to tackle hallucination issues, these often require costly interventions such as fine-tuning with annotated data (Li et al., 2023c), incorporating auxiliary models, or leveraging external knowledge sources.
@@ 2.1. MLLM Development and Capacity
In recent years, Multimodal Large Language Models (MLLMs)(Achiam et al., 2023; Liu et al., 2024; Bai et al., 2023; Zhu et al., 2023; Dai et al., 2024; Li et al., 2023d; Wang et al., 2023b; Team et al., 2023) have rapidly emerged and become the focus of both academia and industry. Since 2021, a series of representative models such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022, 2023b) have pioneered large-scale pre-trained multimodal models, demonstrating the powerful ability to deeply integrate natural language with visual information, enabling accurate image description, cross-modal reasoning, and other functions. Entering 2023, this field is even more explosive, including but not limited to GPT-4V (Achiam et al., 2023), LLaVA (Liu et al., 2024), minGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2024), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and many other new multimodal macromodels have emerged one after another, which further enhance the model’s ability of understanding and generating multimodal inputs, and make the MLLM get closer and closer to the general-purpose Artificial Intelligence (Gutierrez et al., 2023; Bieger et al., 2016) in its ideal form.
@@ 2.3. Strategies for Solving the MLLM Hallucination Problem
The term ”hallucination” (Ji et al., 2023; Zhang et al., 2023b; Montagnese et al., 2021; Gunjal et al., 2024; Li et al., 2023a; Liu et al., 2023b) refers to the fact that multimodal models, when processing multimodal inputs, sometimes produce content that does not correspond to the actual inputs or is even fictitious. Aiming at this key problem, which seriously affects the credibility and practicality of models, the academic community has carried out a lot of fruitful research work and proposed several innovative solutions. Among them, RLHF (Yu et al., 2023; Sun et al., 2023; Zhao et al., 2023b) (Reinforcement Learning from Human Feedback) is an approach that relies on human feedback reinforcement learning techniques, which manually evaluates and guides model outputs, prompting the model to pay more attention to factual basis and logical consistency in the subsequent generation process. DoLa (Chuang et al., 2023) improves the model’s ability to capture and reproduce factual information when generating text by comparing the decoded information at different levels within the model. The Woodpecker (Yin et al., 2023) takes a two-stage approach to hallucination, first optimising the model initially with DINO (Liu et al., 2023c) (a self-supervised learning framework), and then correcting the model’s multimodal outputs by combining it with image captioning techniques. The OPERA (Huang et al., 2023b) research team proposes a novel strategy to effectively mitigate the false inferences and hallucination representations generated by multimodal large models due to over-trusting one modality when processing complex scenes by applying the Over-Trust Penalty and Retrospection Allocation mechanisms. In the specific application scenarios of visual-linguistic modelling, the VCD (Leng et al., 2023) technique helps to identify and correct the object hallucination that may occur when the model is describing or reasoning about an image by introducing the visual contrast decoding link. In conclusion, whether following the traditional route of increasing model size or seeking the innovative path of efficient small-scale models, suppressing the ”hallucination” phenomenon of large multimodal models has become an important issue of common concern to researchers, and substantial progress has been made in several cutting-edge researches. These strategies not only enrich the design and optimisation of multimodal models, but also lay a solid foundation for the construction of more accurate, reliable and universal multimodal intelligent systems in the future.
@@ 3.4. High-Response Region Visualization and Cross-modal Interaction
To explore the relationship between generated text tokens and their correspondence with image attention in a more vivid manner, we visualize the high-response regions of the top 50 scores in Fig. 5. Given user input, we generate a text-guided query vector $Q_{t}\in\mathbb{R}^{M\times C}$, where $M$ denotes the number of queries. As shown in Fig. 3, this cross-modal interaction primarily occurs within the text decoder and can be readily instantiated using BERT (Devlin et al., 2018) or Q-former (Li et al., 2023b) models. The generated text query $Q_{t}$ encapsulates salient visual cues that are most relevant to the user’s command.
@@ 4.2. Quantitative Results
DOPRA evaluation on hallucinations using POPE. The Polling-based Object Probing Evaluation (POPE) (Li et al., 2023a) is a recently developed method aimed at assessing hallucination issues in MLLMs. Analogous to CHAIR, POPE scrutinizes object hallucination by prompting the model with an inquiry akin to ”Is There a ¡object¿ in the image?” to gauge whether the model correctly identifies the presence or absence of a specific object in the given image. POPE includes three distinct evaluation scenarios: ”random”, ”popular”, and ”adversarial”.
@@ 6. Conlusion
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.

### v2

@@ Abstract.
* represents the co-first author, Xiaofeng Zhang is the corresponding author, his email:framebreak@sjtu.edu.cn
copyright: acmlicensedjournalyear: 2024doi: 10.1145/3664647.3681076copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiaccs: Computing methodologies Computer vision tasks
@@ 1. Introduction
Recently, Multimodal Large Language Models (MLLMs)(Achiam et al., 2023; Liu et al., 2024; Bai et al., 2023; Zhu et al., 2023; Zhang et al., 2024; Dai et al., 2024; Li et al., 2023e; Team et al., 2023; Gong et al., 2023; Zhang et al., 2023a) have made groundbreaking advancements, fundamentally altering the way AI interacts with visual data and significantly enhancing fluent communication based on image semantic content. Despite their remarkable performance in handling a range of visually-centered tasks (Zhang et al., 2023c; Black et al., 2023; Li et al., 2024b; Chen et al., 2019; Cai et al., 2023), understanding complex contexts(Lai et al., 2023; Zhao et al., 2023a), or generating coherent narratives (Brooks et al., 2023; Black et al., 2023; Li et al., 2024b), MLLMs still grapple with a profound challenge: the ”hallucination” problem. This refers to instances where MLLMs generate inaccurate or disjointed responses to visual inputs by incorrectly identifying nonexistent objects, attributes, or relationships within provided images. Such errors carry significant risks, particularly in high-stakes applications like autonomous driving (Muhammad et al., 2020; Shafaei et al., 2018), where misinterpreting visual cues could lead to life-threatening situations. While numerous methods (Liu et al., 2023a; Park et al., 2024; Yin et al., 2023; Wang et al., 2024; Zhou et al., 2023) have been proposed to tackle hallucination issues, these often require costly interventions such as fine-tuning with annotated data (Li et al., 2023d), incorporating auxiliary models, or leveraging external knowledge sources.
@@ 2.1. MLLM Development and Capacity
In recent years, Multimodal Large Language Models (MLLMs)(Achiam et al., 2023; Liu et al., 2024; Bai et al., 2023; Zhu et al., 2023; Dai et al., 2024; Li et al., 2023e; Wang et al., 2023b; Team et al., 2023; Li et al., 2024a; Chen et al., 2023b; Li et al., 2023a; Chen et al., 2023a) have rapidly emerged and become the focus of both academia and industry. Since 2021, a series of representative models such as CLIP (Radford et al., 2021) and BLIP (Li et al., 2022, 2023c) have pioneered large-scale pre-trained multimodal models, demonstrating the powerful ability to deeply integrate natural language with visual information, enabling accurate image description, cross-modal reasoning, and other functions. Entering 2023, this field is even more explosive, including but not limited to GPT-4V (Achiam et al., 2023), LLaVA (Liu et al., 2024), minGPT-4 (Zhu et al., 2023), InstructBLIP (Dai et al., 2024), Qwen-VL (Bai et al., 2023), CogVLM (Wang et al., 2023b), and many other new multimodal macromodels have emerged one after another, which further enhance the model’s ability of understanding and generating multimodal inputs, and make the MLLM get closer and closer to the general-purpose Artificial Intelligence (Gutierrez et al., 2023; Bieger et al., 2016) in its ideal form.
@@ 2.3. Strategies for Solving the MLLM Hallucination Problem
The term ”hallucination” (Ji et al., 2023; Zhang et al., 2023b; Montagnese et al., 2021; Gunjal et al., 2024; Li et al., 2023b; Liu et al., 2023b) refers to the fact that multimodal models, when processing multimodal inputs, sometimes produce content that does not correspond to the actual inputs or is even fictitious. Aiming at this key problem, which seriously affects the credibility and practicality of models, the academic community has carried out a lot of fruitful research work and proposed several innovative solutions. Among them, RLHF (Yu et al., 2023; Sun et al., 2023; Zhao et al., 2023b) (Reinforcement Learning from Human Feedback) is an approach that relies on human feedback reinforcement learning techniques, which manually evaluates and guides model outputs, prompting the model to pay more attention to factual basis and logical consistency in the subsequent generation process. DoLa (Chuang et al., 2023) improves the model’s ability to capture and reproduce factual information when generating text by comparing the decoded information at different levels within the model. The Woodpecker (Yin et al., 2023) takes a two-stage approach to hallucination, first optimising the model initially with DINO (Liu et al., 2023c) (a self-supervised learning framework), and then correcting the model’s multimodal outputs by combining it with image captioning techniques. The OPERA (Huang et al., 2023b) research team proposes a novel strategy to effectively mitigate the false inferences and hallucination representations generated by multimodal large models due to over-trusting one modality when processing complex scenes by applying the Over-Trust Penalty and Retrospection Allocation mechanisms. In the specific application scenarios of visual-linguistic modelling, the VCD (Leng et al., 2023) technique helps to identify and correct the object hallucination that may occur when the model is describing or reasoning about an image by introducing the visual contrast decoding link. In conclusion, whether following the traditional route of increasing model size or seeking the innovative path of efficient small-scale models, suppressing the ”hallucination” phenomenon of large multimodal models has become an important issue of common concern to researchers, and substantial progress has been made in several cutting-edge researches. These strategies not only enrich the design and optimisation of multimodal models, but also lay a solid foundation for the construction of more accurate, reliable and universal multimodal intelligent systems in the future.
@@ 3.4. High-Response Region Visualization and Cross-modal Interaction
To explore the relationship between generated text tokens and their correspondence with image attention in a more vivid manner, we visualize the high-response regions of the top 50 scores in Fig. 5. Given user input, we generate a text-guided query vector $Q_{t}\in\mathbb{R}^{M\times C}$, where $M$ denotes the number of queries. As shown in Fig. 3, this cross-modal interaction primarily occurs within the text decoder and can be readily instantiated using BERT (Devlin et al., 2018) or Q-former (Li et al., 2023c) models. The generated text query $Q_{t}$ encapsulates salient visual cues that are most relevant to the user’s command.
@@ 4.2. Quantitative Results
DOPRA evaluation on hallucinations using POPE. The Polling-based Object Probing Evaluation (POPE) (Li et al., 2023b) is a recently developed method aimed at assessing hallucination issues in MLLMs. Analogous to CHAIR, POPE scrutinizes object hallucination by prompting the model with an inquiry akin to ”Is There a ¡object¿ in the image?” to gauge whether the model correctly identifies the presence or absence of a specific object in the given image. POPE includes three distinct evaluation scenarios: ”random”, ”popular”, and ”adversarial”.